OpenStudy (anonymous):

(sq rt of 3 ) / (sq rt of 4)

OpenStudy (mertsj):

\[\frac{\sqrt{3}}{2}\]

myininaya (myininaya):

\[\sqrt{4}=2\]

OpenStudy (anonymous):

square root of 3/square root of 4 Square root of 4 =2 So, final answer: square root of 3/2